### v3

# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

Abstract
CLIP is a foundational multimodal model that aligns image and text features into a shared space using contrastive learning on large-scale image-text pairs. Its strength lies in leveraging natural language as a rich supervisory signal. With the rapid progress of large language models (LLMs), we explore their potential to further enhance CLIP’s multimodal representation learning. This work introduces a fine-tuning approach that integrates LLMs with the pretrained CLIP visual encoder, leveraging LLMs’ advanced text understanding and open-world knowledge to improve CLIP’s ability to process long and complex captions. To address the challenge of LLMs’ autoregressive nature, we propose a caption-to-caption contrastive learning framework to enhance the discriminative power of their outputs. Our method achieves substantial performance gains on various downstream tasks, demonstrating the effectiveness of combining LLMs with CLIP for enhanced multimodal learning.
1 Introduction
CLIP [32] is one of the most influential multimodal foundational models today. It aligns visual and language signals into a shared feature space by employing a simple contrastive learning loss on large-scale image-text pairs. As a retriever, CLIP supports a wide range of tasks, including zero-shot classification [31], detection [22], segmentation [46], and image-text retrieval [27, 17]. As a feature extractor, it has become dominant in virtually all cross-modal representation tasks, such as image understanding, video understanding, and text-to-image/video generation. For instance, works like LLaVA [24] and Qwen-VL [2] leverage CLIP as a feature extractor to obtain visual features for text models, while models like Stable Diffusion [34] and DALL·E 2 [33] use CLIP’s text encoder to extract textual features for visual models.
Figure 1: Zero-shot multimodal retrieval radar chart. LLM2CLIP significantly improves the performance of original models like CLIP and EVA02 through fine-tuning, substantially outperforming other fine-tuning-based methods such as LongCLIP.
Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
Figure 3: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
What makes CLIP particularly powerful as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 [1] and Llama [8] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning. This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
The potential benefits of incorporating LLMs into CLIP are clear. LLMs’ strong textual understanding can fundamentally improve CLIP’s ability to handle image captions, drastically enhancing its ability to process long and complex texts—a well-known limitation of vanilla CLIP. Moreover, LLMs are trained on a vast corpus of text, possessing open-world knowledge. This allows them to expand on caption information during training, increasing the efficiency of the learning process.
In this work, we aim to design a method that incorporates the powerful capabilities of LLMs to fine-tune pretrained state-of-the-art (SOTA) CLIP models, further enhancing their performance. We propose replacing the weak text encoder of CLIP with an LLM, followed by cross-modal contrastive learning training on image-text datasets. While LLMs have already been introduced into CLIP in the InternVL [6] framework, this approach requires large-scale retraining, and the actual performance improvement solely attributable to the LLM remains unclear. In our experiments, we found that directly using LLMs to fine-tune CLIP led to failure in all tasks, with a significant drop in performance across various downstream tasks as in Table 6. We discovered that the output features of LLMs for different texts (or even the same text) were very similar, which led to poor differentiation. This low discriminative power interferes with CLIP’s contrastive learning task. To investigate this further, we designed a caption-to-caption retrieval experiment, as shown in Table 1 and Figure 4. In this experiment, each image in the MS-COCO dataset has five human-annotated captions. We selected the first two captions as positive samples and performed retrieval across the entire test set. Using caption retrieval accuracy (CRA), we evaluated the text model’s ability to differentiate between captions, helping us determine which language model is better suited for CLIP. We found that Llama-3 8B achieved only 5.2% top-1 accuracy, while the standard CLIP-ViT-L reached 25.2% top-1 accuracy. As illustrated in Figure 4, the top-1 caption retrieved by the original Llama-3 model can be entirely unrelated to the query caption, clearly obstructing effective CLIP learning. Therefore, directly using an LLM to guide CLIP’s visual encoder training is highly constrained. Recent work on LLM2Vec also validates this idea, leading us to realize the necessity of fine-tuning the output space of LLMs.
Language Model CRA
CLIP-L/14 25.2
EVA02-L/14 27.1
Llama3-8B 5.2
Llama3.2-1B 5.6
Llama3-8B-CC 29.5
Llama3.2-1B-CC 29.4
Table 1: Comparison of top-1 Caption Retrieval Accuracy (CRA) for various language models on the MS COCO 5K test set.
Figure 4: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
We designed a caption contractive (CC) fine-tuning strategy, as Figure 2, pplying lightweight fine-tuning to the output tokens of Llama-3 8B using LoRA on the CC3M [36] image captioning dataset. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE [11, 3] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [5] were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%. This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP. As in  Figure 3, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining. LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space. In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance. Experimental results demonstrate that training on only 15M open-source data substantially boosts CLIP’s capabilities, with even 3M data yielding impressive results.
The main contributions of this paper are:
  1. 1.
    We identify the core limitation preventing LLMs from fully leveraging their capabilities within the CLIP training framework—the insufficient separability of LLM output features.
  2. 2.
    We introduce a straightforward approach by applying caption-to-caption contrastive fine-tuning on the LLM’s output layer, enhancing the separability of the output space. This approach shows experimentally that involving LLMs in the CLIP training process can provide substantial benefits.
  3. 3.
    We propose LLM2CLIP, an efficient and effective LLM-assisted fine-tuning method for CLIP, enabling low-cost fine-tuning of pretrained CLIP models and achieving significant performance improvements.
2 Related Works
CLIP meets Stronger Language Models.
Several works have explored the integration of LLMs into CLIP. JinaCLIP [17] employed Jina-embeddings-v2 [13] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP [38], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. E5-V [15] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [19] while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE [14] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
CLIP meets Longer Captions.
It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI [40] hired annotators to expand captions based on SAM [16]’s target hints. LaCLIP [9] adopted ChatGPT, Bard, and human rewriting. DreamLIP [45] leveraged ShareCaptioner [5] and InstructBLIP to augment 30M captions. Recap-DataComp-1B [20] used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [40], splitting captions into multiple segments [9, 45], or finetuning model’s positional encoding to support longer token inputs [44]. LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
3 Methods
LLM2CLIP is a fine-tuning strategy for CLIP, consisting of two stages. The first stage aims to improve the separability of the LLM’s output space, and the second stage involves fine-tuning by replacing CLIP’s text model with the LLM.
3.1 Caption Contrastive Fine-tuning
In this section, we aim to fine-tune the LLM’s token outputs to better capture features that can distinguish between image captions.
The process of improving the discriminability of LLM output features on caption text is quite straightforward: we need the distances between captions of the same image to be closer, and those of different images to be further apart. Therefore, we apply a caption contrastive (CC) fine-tuning to the LLM’s output features, treating different captions of the same image as positive samples and the rest of the captions as negative samples.
To obtain enough varied descriptions for the same image, we use the ShareCaptioner [45, 5] modified CC-3M [36] dataset, which provides both original captions and augmented dense captions for each image. These can be treated as positive pairs. We followed the training methodology of LLM2Vec [3], first expanding the LLM’s attention mechanism to bidirectional attention, and employing Masked Next Token Prediction (MNTP) for initialization to achieve better results. Specifically:
  1. 1.
    Bidirectional Attention Transformation: First, we transform the LLM’s causal attention mechanism into bidirectional attention, as we no longer need it to retain generative capabilities but rather function as an encoder. Since autoregressive training is no longer required, switching to bidirectional attention improves its ability to capture contextual information.
  2. 2.
    Masked Next Token Prediction (MNTP): Second, we employ MNTP to train the newly added bidirectional attention mechanism, providing a strong initialization. For a given sequence of $N$ tokens, we mask a subset and predict their values, similar to BERT [7]. However, unlike BERT, we adapt this process to fit the nature of LLMs by predicting tokens just before the masked token. We train on both image captions and pure text with equal weighting. In addition to CC-3M, we also use the Wikitext-103 [28] dataset to preserve the LLM’s text capabilities and prevent divergence from its original strengths.
  3. 3.
    Caption Contrastive Learning: Finally, we perform the actual caption contrastive fine-tuning, using a supervised SimCSE loss to bring captions of the same image closer together and push apart captions of different images. We use two prompt templates: “Given a caption, retrieve a detailed relevant caption” and “Given a detailed caption, retrieve a short relevant caption,” which are prepended to the query (either original or dense caption) to retrieve the corresponding dense or original caption. Similarly, we use a 1.5M-paired general text dataset curated from [37] to maintain strong performance in pure language tasks.
All training is efficiently conducted using LoRA and is completed in just one epoch, ensuring low computational cost. In our experiments, we found that the most critical step was the third step with the SimCSE loss, which significantly enhanced the caption discriminability of the LLM.
3.2 LLM2CLIP: Use LLM to Finetuning
With the modifications to the LLM discussed earlier, we have now obtained a super text encoder that is well-suited for CLIP training. The next step is to use this LLM in conjunction with the pretrained state-of-the-art CLIP visual encoder to reconstruct a more powerful cross-modal feature space.
As shown in  Figure 3, in the LLM2CLIP training phase, we freeze the gradients of the LLM to preserve its inherent capabilities, for two primary reasons. First, this significantly reduces the computational cost and memory footprint of fine-tuning. CLIP training requires a very large batch size to maintain the effectiveness of negative samples. Allocating memory to the LLM could compromise CLIP’s performance. Second, by freezing the LLM, we ensure that the open-world knowledge it has acquired from large-scale corpora remains intact during the multimodal alignment process.
To compensate for the frozen LLM, and inspired by methods like FuseMix [41] and APE [35], we introduce several new linear layers after the LLM as adapters. These layers serve as learnable parameters to improve the alignment between the LLM and CLIP visual encoder. Following the original design of CLIP, we also employ a projector layer to align the dimensions of the two encoders, facilitating the use of CLIP loss for training.
With this powerful LLM-based super text encoder, we achieve a substantial qualitative leap in CLIP’s language comprehension capabilities as results in Table 2. The LLM’s open-world knowledge allows the CLIP visual encoder to learn more structured and globally informed visual representations that are aligned with human knowledge. Moreover, this approach enables us to fully utilize high-quality, long, and dense caption datasets without requiring any special architectural adjustments, which previous works like DCI, DreamLip, and Recaption struggled to effectively leverage.
3.3 Overview and Efficiency Discussion
We propose LLM2CLIP as a method that efficiently incorporates large language models (LLMs) into CLIP training, leveraging the capabilities of LLMs to make cross-modal representation learning significantly more powerful. In our experiments, We evaluated large language models, including Llama with 1B and 8B parameters, as well as Mistral-Nemo with 12B parameters.It might seem that incorporating such large LLMs would greatly increase the computational burden of training CLIP, especially since CLIP itself is computationally expensive, requiring a large batch size. However, our proposed LLM2CLIP is remarkably lightweight. The training overhead isnearly identical to fine-tuning the original CLIP model, with minimal additional cost, yet the LLM provides much stronger supervision.
Here, we highlight some of the design details that significantly improve training efficiency: 1). During the caption contrastive fine-tuning stage, we employ LoRA training for the LLM. Even for a 12B LLM, training with a batch size of 512 only requires around 70GB of GPU memory, making it feasible to run on a single 80GB 8 A100 GPU node. 2). In the LLM2CLIP stage, we freeze the LLM’s gradients and only train the learnable adapter, CLIP’s original Vision encoder, and two projectors. The additional trainable parameters are roughly equivalent to those in the original CLIP, minimizing the overhead. To further reduce the inference cost of using LLMs, we pre-extract all the text features from the training data and store them in disk memory. This ensures that, even though LLMs provide powerful textual supervision, the memory and computational costs during training remain nearly identical to those of standard CLIP training.
For instance, when we trained LLM2CLIP using a Mistral-Nemo [^1] https://mistral.ai/news/mistral-nemo/ 12B model integrated with the commonly used EVA ViT-L/14-224, with a batch size of 4096 on 8 H100 GPUs, the memory usage per GPU was only 30GB, and the entire training process took just 9 hours. Despite this efficient training cost, LLM2CLIP brought transformative improvements in downstream tasks such as long and short text retrieval, cross-language retrieval, and Llava training.
Methods Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
ViT-B/16
ALIGN 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Jina-CLIP 80.6 67.4 55.6 41.1 - - 87.7 88.0 78.7 80.0 75.7 69.1
Long-CLIP 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
CLIP 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
+LLM2CLIP 89.2 78.1 62.2 48.7 98.1 97.4 86.1 90.0 84.1 85.0 83.9 79.8
EVA02 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
+LLM2CLIP 88.5 78.0 63.6 49.8 98.0 98.1 84.7 89.7 85.5 86.8 84.1 80.5
ViT-L/14
Long-CLIP 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
CLIP 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
+LLM2CLIP 92.6 81.7 64.9 52.5 98.4 98.4 87.6 92.0 87.6 88.7 86.2 82.7
EVA02 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+LLM2CLIP-3M 89.6 77.3 59.7 48.0 98.3 98.6 87.1 91.1 84.9 87.8 83.9 80.6
+LLM2CLIP 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
+LLM2CLIP-30M 92.0 83.5 69.0 55.3 98.9 98.8 93.1 95.0 89.3 91.2 88.5 84.8
+LLM2CLIP-60M 94.4 83.2 70.4 55.7 99.2 99.4 94.1 95.2 90.2 92.0 89.7 85.1
ViT-L/14-336
CLIP 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
+LLM2CLIP 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
+LLM2CLIP-60M 93.9 82.3 68.5 54.8 98.9 99.1 94.6 95.9 89.6 90.6 89.1 84.5
EVA02 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
+LLM2CLIP 93.9 83.8 68.7 55.7 98.8 99.2 89.5 94.2 89.2 91.3 88.0 84.8
Table 2: Systematic Comparison Experiment Demonstrating the Performance Improvements of LLM2CLIP. Results with a blue background represent the outcomes of LLM2CLIP.
4 EXPERIMENTS
4.1 Experiment Setting
4.1.1 Caption Contrastive Fine-tuning
Training Dataset.
For contrastive learning, we utilized the ShareCaptioner-modified CC-3M dataset [45, 5], which includes both original captions and augmented dense captions for each image. In addition, we used the Wikitext-103 dataset [28] and the E5 dataset [37] during the Masked Next Token Prediction and caption contrastive fine-tuning stages to ensure coverage of the general text domain without excessive bias.
Training Setting.
The language model was trained with LoRA for lightweight training, completing one epoch across all datasets. We used the average of all LLM output tokens as the global text embedding for a caption. All training parameters followed the configuration detailed in [3]. The design of the Adaptor is the same as FuseMix [41], using a simple inverted bottleneck MLP architecture.
4.1.2 LLM2CLIP Fine-tuning
Training Dataset.
We defined three experimental configurations based on dataset size:
  • LLM2CLIP-3M: This configuration uses only the CC-3M dataset, serving as our lightweight setting.
  • LLM2CLIP-15M: Our default configuration, combining CC-3M and CC-12M datasets.
  • LLM2CLIP-60M: A larger configuration that combines CC-3M, CC-12M, YFCC-15M, and a 30M randomly selected subset from Recaption-1B [20].
All datasets included dense captions rewritten by multimodal large language models (MLLMs). Specifically, the CC-3M, CC-12M, and YFCC datasets were sourced from [45] and rewritten using ShareCaptioner, while Recaption data were rewritten with Llama3-LLAVA1.5. A 1:1 mix of original captions and dense captions was used by default.
Training Setting.
Each configuration was trained for four epochs. We froze the LLM gradients to reduce computational overhead and pre-extracted text features using the LLM, storing them on disk to avoid repeated inference during training. Inspired by Fusemix’s efficient training strategy, we added a four-layer linear adapter after the LLM and trained with gradients enabled for the CLIP ViT model to better integrate LLM knowledge. Additionally, we introduced two projector layers after each encoder for dimensional alignment.
Training the CLIP ViT-L/14-224 model required approximately 30GB of GPU memory per GPU and took 2, 9, and 45 hours on 8 H100 GPUs for the 3M, 15M, and 60M datasets, respectively. We employed the AdamW [26] optimizer, utilizing a total batch size of 4096, and set the learning rate to 1e-5 with a cosine decay rate of 0.05.
For language models, we tested four configurations: Jina-Embeddings-V2, and three popular models from the Llama3 family [8]: Llama 3.2 1B, Llama 3 8B, and Mistral-Nemo 12B. Unless otherwise noted, experiments used EVA02 ViT-L/14 as the baseline and Llama 3 8B as the default language model.
Methods Flickr-CN COCO-CN
I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10 I2T@1 I2T@5 I2T@10 T2I@1 T2I@5 T2I@10
ViT-L/14-336
Wukong 76.1 94.8 97.5 51.7 78.9 86.3 53.4 80.2 90.1 55.2 81.0 90.6
CN-CLIP 80.2 96.6 98.2 68.0 90.7 95.4 63.4 84.2 92.9 64.0 89.2 94.4
JinaCLIP 3.30 9.90 15.1 0.7 3.5 6.0 2.9 8.9 13.7 1.0 4.9 8.2
EVA02 4.40 11.8 16.7 0.94 2.9 4.8 2.7 9.8 15.2 1.0 3.7 7.3
+LLM2CLIP 86.9 98.1 99.3 75.1 92.9 96.0 69.1 92.5 97.2 70.0 92.6 96.7
Table 3: Retrieval Performance across Flickr30K-CN and COCO-CN.
Methods Imagenet
0-shot Linear
CLIP L/14-336 76.6 84.8
+ LLM2CLIP 74.1 85.2
CLIP L/14 75.5 83.9
+ LLM2CLIP 71.8 84.4
CLIP B16 70.2 79.9
+ LLM2CLIP 67.3 80.6
Table 4: Imagenet 0-shot classification and linear probe results.
MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
Llava (Paper) 78.5 62.0 50.0 66.8 58.2 87.3 86.1 84.2 1510.7 64.3 58.3 65.4 31.1 58.6 66.1 37.3
Llava (Rep.) 79.04 62.86 50.57 67.97 57.48 87.7 84.85 86.3 1476.69 66.66 60.39 58.0 34.3 59.86 66.95 39.71
+LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
Table 5: Performance of Llava 1.5 benchmarks. The best results are highlighted in bold.We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average CRA
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5 27.1
+Jina Bert 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0 29.68
+Llama3-8B 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0 5.22
+Llama3-8B-TC(LLM2Vec) 92.7 82.1 68.1 54.6 97.7 98.2 88.9 93.8 85.0 87.8 86.5 83.3 28.42
+Llama3-8B-CC 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2 29.46
+Llama3.2-1B-CC 91.6 81.3 65.8 52.5 98.3 98.2 84.5 91.9 83.4 86.4 84.7 82.1 29.42
+Mistral-Nemo-12B-CC 93.5 83.7 68.5 54.7 98.6 98.9 90.4 94.3 88.0 89.7 87.8 84.3 29.96
Table 6: Comparison of various text encoders.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
EVA02 Vit-L/14 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
+ Jina-Bert 88.1 77.7 60.5 51.1 83.3 81.0 66.9 68.5 68.9 71.2 73.5 70.0
       ++ Dense Caption 87.9 77.9 60.9 50.3 95.3 95.1 79.4 83.8 73.8 77.9 79.5 77.0
+ Llama3-8B 87.9 75.6 56.7 41.8 55.1 46.1 37.2 35.1 39.3 32.3 55.2 46.2
       ++ Dense Caption 87.1 75.3 56.4 41.6 89.3 91.4 58.6 60.9 51.7 50.6 68.6 64.0
+ Llama3-8B-CC 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
       ++ Dense Caption 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
Table 7: Ablation analysis of performance variations between short simple captions and long & dense captions using different text encoders. We conducted experiments on the EVA02 ViT-L/14 model, replacing its original text encoder with Jina-BERT (+ Jina-BERT), the original Llama3 (+ Llama3-8B), and Llama3 fine-tuned with Caption Contrastive learning (+ Llama3-8B-CC). Fine-tuning was performed on datasets containing only short simple captions and datasets with a 50% probability mix of long dense captions. The objective of this experiment was to isolate the effects of LLMs and long dense captions, and to demonstrate that the design of LLM2CLIP enhances the efficiency of dense caption utilization. The results with a red background indicate that the LLM without fine-tuning has a significant negative effect on CLIP.
Ratio Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
100% 85.5 72.7 60.1 46.9 98.7 99.0 88.7 93.9 88.0 90.5 84.2 80.6
75% 92.4 82.6 68.5 54.2 98.7 99.3 89.0 94.3 88.1 90.2 87.3 84.1
50% 92.0 82.8 68.5 54.8 98.6 99.0 88.1 94.0 88.2 90.4 87.1 84.2
25% 93.0 82.8 68.1 54.8 98.4 98.7 87.7 92.9 87.9 90.0 87.0 83.8
0% 92.4 82.9 67.6 54.5 97.7 94.9 75.8 83.4 83.7 85.6 83.4 80.3
Table 8: Comparison Experiment of Different Ratios of Dense Captions in the LLM2CLIP Training Process.
Adaptor Layers Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
0 91.3 79.9 64.0 50.2 97.1 97.7 88.1 90.2 81.0 86.1 84.3 80.8
1 92.1 80.6 65.9 51.8 97.8 98.0 90.9 91.8 84.3 87.9 86.2 82.0
4 92.1 79.9 66.0 52.2 97.4 98.0 93.0 92.1 85.0 88.0 86.7 82.0
Table 9: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different adaptor layer numbers. Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks. The best result in each column is highlighted in bold. All experiments were conducted using OpenAI’s CLIP model with a ViT architecture at 224 resolution, trained on 15M data points.
Methods Flickr30k COCO ShareGPT4v Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
Bidirectional Attention + MNTP + Sup. SimCLR 83.3 73.5 55.3 45.6 98.1 98.0 81.5 87.9 78.8 82.6 79.4 77.5
Bidirectional Attention + Sup. SimCLR 84.6 74.5 55.6 46.2 98.0 97.6 79.3 88.5 80.0 82.6 79.5 77.9
Sup. SimCLR 82.0 73.5 54.6 44.8 97.6 97.6 82.2 87.9 76.9 81.5 78.7 77.1
Table 10: Performance comparison across benchmarks (Flickr30k, COCO, ShareGPT4v, Urban-1k, and DOCCI) using different methods. Results are reported for image-to-text (I2T) and text-to-image (T2I) retrieval tasks. The best result in each column is highlighted in bold. The experiments were conducted using the Llama 3.2 1B and the EVA02 ViT-L/14 model on 3M data. It can be observed that MNTP has no significant effect, while bidirectional attention provides a slight performance improvement.
4.2 Main Results
Evaluation Setting.
Our experiments assessed LLM2CLIP’s ability to enhance vanilla CLIP performance across multiple dimensions. We compared our approach against state-of-the-art (SOTA) models, including EVA [10] and OpenAI CLIP. Specifically, we evaluated CLIP ViT-B/16, ViT-L/14, and ViT-L/14-336, along with EVA02 ViT-B/16, EVA02 ViT-L/14, and EVA02 ViT-L/14-336, yielding a total of six comparison models.
4.2.1 Short- and Long-Text Retrieval
Evaluation Dataset.
For short-text retrieval, we used the MSCOCO [23] 5K test set and the Flickr [43] 1K test set. For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [5], the Urban1K [44] dataset, and the DOCCI [29] dataset.
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [30], and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval.
As shown in Table 2, LLM2CLIP fine-tuning yielded significant performance improvements, even surpassing models pretrained on large datasets like LAION2B. Compared to LongCLIP and JinaCLIP, our method achieved transformative gains, showcasing its efficiency in leveraging LLM knowledge to enhance pretrained CLIP models.
4.2.2 Cross-Lingual Retrieval
Evaluation Dataset.
For Chinese retrieval tasks, we used FlickrCN [18] and CNCOCO [21], which are translated versions of Flickr30K and MS COCO-1K, respectively. LLAVA training and evaluation were conducted using the standard splits from LLAVA 1.5 [24].
LLM Enables English CLIP to Learn Chinese Without Exposure to Chinese Data.
To test LLM knowledge transfer, we conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. As shown in Table 3, EVA02 CLIP and JinaCLIP showed near-zero performance, while LLM2CLIP achieved remarkable results, even outperforming models trained on extensive Chinese data like Wukong [12] and CN-CLIP [42]. This demonstrates the power of LLMs in transferring knowledge to cross-modal spaces.
4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
Evaluation Dataset.
We conduct zero-shot retrieval testing on Imagenet, following the CLIP method. We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
Imagenet Performance Better In Linear Probe.
Our method shows a slight drop in performance on Imagenet for zero-shot classification. We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions. However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement. This demonstrates that LLM2CLIP can indeed provide better visual features. One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
4.2.4 Multimodal Large Language Models Performance
Evaluation Dataset.
Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
LLM2CLIP Improves VLLM Performance.
As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
4.3 Main Conclusions
Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
Directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance, as shown by the red-highlighted numbers in Table 7 and Table 6, with average performance drops of 10%-20%. Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features lack sufficient discriminability for captions. These results clearly demonstrate that such an approach imposes a significant burden on CLIP’s learning process.
Improving the Discriminability of LLM Output Features is Key to Integrating LLMs with CLIP.
To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning. Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora(same as LLM2Vec) and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss. As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks. This difference stems from the fact that LLMs trained on caption-distributed data have stronger discriminability for caption features. These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
LLM’s Built-in Knowledge Directly Benefits CLIP.
In Table 7, we observe that directly using LLM2CLIP leads to significant performance improvements even with short caption datasets, which is quite challenging, as CLIP has already seen most of this training data. However, the results from LLM2CLIP with +Llama3-8B-CC still show a clear performance gain, which can be attributed purely to the improvement in the text encoder. On the other hand, +Jina-Bert fails to achieve similar results, and in fact, leads to a performance decline.
LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications. Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT. To verify the LLM’s natural advantage with long texts, we compared our method to Jina CLIP to assess whether the LLM provides a better understanding of long captions. As shown in Table 7, when dense captions are added during training, LLM2CLIP demonstrates significantly greater performance improvements than Jina CLIP, proving that LLMs are better suited to processing long and dense captions.
Larger LLMs Lead to Better Performance.
LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training. We thoroughly investigate the performance differences that various versions of LLMs bring to CLIP. In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact. We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder. As expected, within the same Llama family, the 8B model significantly outperfoWrms the 1B model. Mistral Memo also shows a slight improvement over the 8B model. It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP.
4.4 Ablation
Impact of Different Data Ratios.
With the integration of an LLM, LLM2CLIP enhances our ability to understand dense captions. In  Table 8, we present an ablation study that examines the impact of different ratios of dense captions versus original captions on CLIP’s training performance. The “Ratio” column represents the proportion of dense captions used during training. Overall, dense captions contribute to a noticeable improvement in performance, but more is not always better. For example, when using 0% dense captions, the performance on datasets like COCO and Flickr30K remains decent, but long-text benchmarks such as ShareGPT4V, Urban-1K, and DOCCI show poor results. Conversely, with 100% dense captions, performance on short-text retrieval benchmarks is the worst. It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal. Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data.
Impact of Different Data Size.
As shown in  Table 2, larger training datasets consistently yield positive results for LLM2CLIP, with clear improvements in both long-text and short-text retrieval tasks. Our 60M dataset version has already pushed the limits of what CLIP could previously achieve. Even the 3M version, though lightweight, still delivers significant performance gains, demonstrating the efficiency of the LLM2CLIP approach. It’s worth noting that training on the 3M dataset takes only about 3 hours on an 8 H100 GPUs machine, yet results in a transformative leap in performance for a well-pretrained CLIP ViT model.
Adaptor design.
In Table 9, we explore the impact of the Adaptor and its layer count. We find that adding at least one layer of the Adaptor is beneficial, resulting in an average performance improvement of two points. However, increasing it to four layers provides minimal additional gains.
Caption Contrastive Fine-tuning Ablation.
In Table 10, we explore the impact of the key steps in LLM2Vec’s Caption Contrastive fine-tuning, namely MNTP and bidirectional attention. Through comparisons, we observe that MNTP offers almost no positive effect, while bidirectional attention provides some performance gains but is not decisive. Therefore, we conclude that the most significant contribution comes from the supervised SimCSE contrastive learning applied to the captions.
5 Limitations and Future Work
This paper focuses solely on the impact of LLMs on CLIP. In reality, there are many other areas where text can play a transformative role, such as diffusion models [34, 4, 39], Segment Anything [16] and GroundingDINO [25], both of which could potentially see significant performance improvements by applying our method.
While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential. LLM2CLIP is a method that requires only a small amount of data for fine-tuning in conjunction with a pretrained CLIP visual encoder, and we have already demonstrated significant performance gains. However, we did not intentionally select fine-tuning data based on specific characteristics. By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks.
6 Conclusion
Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications. The core contributions of this work are threefold: First, we identified the primary reason why LLMs struggle to directly participate in multimodal representation learning — the lack of discriminability in their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability, removing the biggest obstacle to utilizing LLMs in CLIP training. Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models.