### v4

# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

Abstract
CLIP is a foundational multimodal model that aligns image and text features into a shared representation space via contrastive learning on large-scale image-text pairs. Its effectiveness primarily stems from the use of natural language as rich supervision. Motivated by the remarkable advancements in large language models (LLMs), this work explores how LLMs’ superior text understanding and extensive open-world knowledge can enhance CLIP’s capability, especially for processing longer and more complex image captions. We propose an efficient post-training strategy that integrates LLMs into pretrained CLIP. To address the challenge posed by the autoregressive nature of LLMs, we introduce a caption-to-caption contrastive fine-tuning framework, significantly enhancing the discriminative quality of LLM outputs. Extensive experiments demonstrate that our approach outperforms LoRA-based methods, achieving nearly fourfold faster training with superior performance. Furthermore, we validate substantial improvements over state-of-the-art models such as CLIP, EVA02, and SigLip2 across various zero-shot multimodal retrieval tasks, cross-lingual retrieval tasks, and multimodal language model pretraining.
1 Introduction
Figure 1: LLM2CLIP Overview. After applying caption contrastive fine-tuning to the LLM, the increased textual discriminability enables more effective CLIP training. We leverage the open-world knowledge and general capabilities of the LLM to better process dense captions, addressing the previous limitations of the pretrained CLIP visual encoder and providing richer, higher-dimensional textual supervision.
Figure 2: By adjusting the output space of the LLM, we enable the LLM to more effectively act as a private tutor for CLIP, allowing the fine-tuning process of LLM2CLIP to efficiently adjust the pretrained CLIP’s cross-modal space.
CLIP [28] is one of the most influential multimodal foundational models today. What makes CLIP particularly powerful as a vision encoder? The core of its strength lies in its unprecedented ability to align visual pretraining with natural language—the carrier of human knowledge. Unlike earlier vision encoders such as VGG and ResNet, which relied on the limited ImageNet dataset and simple image categories with just a few words, CLIP is trained on web-scale data using rich descriptive text. This alignment with language is what sets CLIP apart and unlocks its vast potential. However, since CLIP’s introduction, large language models (LLMs) have advanced significantly. Models like GPT-4 [1] and Llama [6] now demonstrate remarkable language capabilities, yet these advancements have not translated into corresponding improvements in visual representation learning. This raises the question: Can the capabilities of LLMs be harnessed to further improve multimodal representation learning?
The potential benefits of incorporating LLMs into CLIP are clear. LLMs’ strong textual understanding can fundamentally improve CLIP’s ability to handle image captions, drastically enhancing its ability to process long and complex texts—a well-known limitation of vanilla CLIP. Moreover, LLMs are trained on a vast corpus of text, possessing open-world knowledge. This allows them to expand on caption information during training, increasing the efficiency of the learning process.
This paper successfully explores a efficient method of leveraging LLM capabilities in CLIP pretraining. LLMs, as text generation models, are primarily pretrained to capture the distribution of natural language, thus their outputs naturally lie close to textual spaces. However, what CLIP pretraining actually requires is a robust text embedding model. Previous works such as LLM2Vec [2] and NV-Embed-v2 [17] have demonstrated that fine-tuning LLMs significantly enhances their ability as text encoders. Nevertheless, it remains unclear how directly applying LLMs affects CLIP performance. Our preliminary results indicate that directly integrating LLMs into CLIP pretraining can be detrimental because LLMs struggle to differentiate effectively between image descriptions, as illustrated by example cases in Figure 3. Moreover, as shown in Table 1, we conducted caption-to-caption retrieval experiments on the COCO 5K test set (each image has five captions; we used the first two captions as positive pairs) and observed that large language models like Llama 3 8B achieve very low accuracy, barely distinguishing between different captions of the same image.
Language Model Top 1 Acc.
CLIP-L/14 25.2
EVA02-L/14 27.1
Llama3-8B 5.2
Llama3.2-1B 5.6
Llama3-8B-CC 29.5
Llama3.2-1B-CC 29.4
Table 1: Comparison of top-1 retrieval accuracy for various language models on the MS COCO [22] 5K test set.
Figure 3: Real examples of top-1 results from the caption-to-caption retrieval experiment in MS COCO 5K test set. Before fine-tuning, Llama3’s results were often completely unrelated.
We designed a caption contractive (CC) fine-tuning strategy, applying lightweight fine-tuning to the output tokens of LLM using image captioning dataset as in  Figure 2. The primary goal of this training task was to adjust the output space, improving the model’s ability to distinguish between different captions. We utilized a supervised SimCSE [8, 2] contrastive learning loss, where the original captions and re-annotated captions generated by ShareCaptioner [3] were treated as positive pairs, pulling them closer. In contrast, all other captions formed a negative sample set that the model learned to push away. Remarkably, after this CC fine-tuning, the caption retrieval accuracy, as shown in Table 1, rose from 5.2% to 29.5%. This successful fine-tuning process enables us to more effectively harness the capabilities of LLMs for CLIP training.
Building on fine-tuning of LLMs, we designed an efficient approach, LLM2CLIP, to enhance the performance of pretrained CLIP. As in  Figure 1, This method uses an LLM to further fine-tune the pretrained CLIP, allowing the LLM to impart knowledge and capabilities to CLIP’s vision encoder that were lacking during its initial pretraining. LLM2CLIP consists of two stages: in the first stage, we apply caption contrastive (CC) fine-tuning on the LLM with a small dataset to adjust the LLM’s output space. In the second stage, we fine-tune the CLIP vision encoder with the CC-fine-tuned LLM, with LLM gradients frozen, to further improve CLIP’s performance.
The main contributions of this paper are:
  1. 1.
    We identify the core limitation preventing LLMs from fully leveraging their capabilities within the CLIP training framework—the insufficient separability of LLM output features.
  2. 2.
    We introduce a straightforward approach by applying caption-to-caption contrastive fine-tuning on the LLM’s output layer, enhancing the separability of the output space. This approach shows experimentally that involving LLMs in the CLIP training process can provide substantial benefits.
  3. 3.
    We successfully verified that LLM2CLIP can substantially enhance the performance of CLIP and EVA02 across multiple tasks, including short-text retrieval, long-text retrieval, cross-language retrieval, and multimodal large language model pretraining, through a simple post-training strategy. Remarkably, LLM2CLIP can further improve the short-text retrieval accuracy of state-of-the-art large-scale pretrained models like SigLip2 [33] by +1 and +1.9 points, achieve impressive gains of +14.8 and +15.8 points in long-text retrieval, and deliver significant improvements of +11.9 and +15.2 points in multilingual tasks.
2 Related Works
CLIP meets Stronger Language Models. Several works have explored the integration of LLMs into CLIP. JinaCLIP [14] employed Jina-embeddings-v2 [9] as the text encoder, which is a BERT variant with 137M parameters, supporting longer texts. Though achieving similar visual performance to EVA-CLIP [30], its text encoder is far behind our used Llama3-8B, limiting the potential benefits from LLMs. E5-V [12] took a different approach by aggregating LLM features at the MLLM layer of LLAVA-next-8B [18] while freezing the ViT gradient to focus on adjusting the MLLM output. However, freezing visual encoder does not address the problem that ViT inherently lack complex visual feature extraction capabilities, leading to much worse performance than LLM2CLIP. MATE [10] designed a learnable adaptor to bridge the gap between CLIP’s text encoder and LLMs. They trained the CLIP visual encoder using LoRA on a small dataset focused on long-text image retrieval tasks. However, they did not recognize the critical issue we propose in this paper: the poor separability of LLM feature space, which is insufficient for direct support of CLIP training. Our work aims to thoroughly explore the impact of LLMs on CLIP as a foundational model, going beyond the influence on long-text retrieval.
CLIP meets Longer Captions. It has been widely recognized that the quality of CLIP’s text embedding is coarse and limited to only 77 tokens. Many works have attempted to extend the length of CLIP captions and retrain CLIP accordingly. For instance, DCI [34] hired annotators to expand captions based on SAM [13]’s target hints. LaCLIP [7] adopted ChatGPT, Bard, and human rewriting. DreamLIP [41] leveraged ShareCaptioner [3] and InstructBLIP to augment 30M captions. Recap-DataComp-1B [21] used Llama3-trained LLAVA1.5 to extend 1B captions. In order to handle captions longer than 77 tokens, these methods compromised by summarizing long captions into shorter ones [34], splitting captions into multiple segments [7, 41], or finetuning model’s positional encoding to support longer token inputs [39]. LLM2CLIP, in comparison, leverages Llama3 as the text encoder, which enables comprehensive understanding of long and dense captions. It not only resolves the limition of token length but also allows for better understanding using LLMs’ open-world knowledge. As a result, our model achieves a breakthrough in performance.
Methods Resolution Flickr30k COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
ViT-B/16(86M)
ALIGN [11] 224 80.6 62.2 52.0 43.2 75.9 80.6 62.2 59.1 59.7 62.1 66.1 61.4
BLIP [19] 224 80.6 74.1 61.7 48.5 65.8 74.3 45.5 48.5 50.5 53.5 60.8 59.8
Long-CLIP [39] 224 85.8 70.6 56.9 40.9 94.8 93.5 79.1 79.1 63.1 71.4 75.9 71.1
jina-clip-v2 [15] 224 84.4 69.8 57.1 41.9 94.5 91.1 80.4 78.0 77.6 78.2 78.8 71.8
SigLIP [38] 224 89.1 74.7 65.8 47.8 85.8 83.5 62.8 62.4 70.1 70.6 74.7 67.8
MetaCLIP [36] 224 85.6 70.8 59.3 41.3 90.4 86.6 68.9 63.3 70.9 71.5 75.0 66.7
CLIP [28] 224 82.3 62.2 52.4 33.1 84.5 79.8 67.5 53.1 60.7 57.1 69.5 57.1
     +LLM2CLIP-15M 224 91.8 80.7 64.8 52.4 97.8 97.6 90.1 92.1 86.1 86.9 86.1 81.9
EVA02 [31] 224 86.2 71.5 58.7 42.1 90.5 85.5 67.0 60.8 67.7 68.0 74.0 65.6
     +LLM2CLIP-15M 224 91.2 81.4 66.0 53.9 98.5 98.5 88.2 91.3 86.8 88.5 86.1 82.7
ViT-L/14 (307M)
Long-CLIP [39] 224 90.0 76.2 62.8 46.3 97.2 97.3 82.5 86.1 66.5 78.6 79.8 76.9
MetaCLIP [36] 224 90.1 76.4 64.5 47.1 86.4 79.5 73.4 70.0 76.5 76.7 78.2 69.9
SAIL-L-NV2 [40] 224 87.9 75.8 62.4 48.7 94.8 94.9 81.5 80.2 76.5 78.9 80.6 75.7
CLIP [28] 224 85.2 65.0 56.3 36.5 84.2 83.6 68.3 55.6 63.1 65.8 71.4 61.3
     +LLM2CLIP-15M 224 93.0 83.5 67.4 56.4 98.7 98.6 91.5 94.1 88.3 90.4 87.8 84.6
EVA02 [31] 224 89.7 77.3 63.7 47.5 91.9 89.3 73.3 68.5 73.5 75.0 78.4 71.5
     +LLM2CLIP-3M 224 94.3 84.3 65.2 56.0 98.0 98.0 91.0 94.7 87.8 90.7 87.3 84.7
     +LLM2CLIP-15M 224 94.6 85.0 69.5 58.3 98.9 99.1 93.6 95.7 89.8 91.2 89.3 85.8
     +LLM2CLIP-60M 224 95.9 85.1 71.7 58.5 99.2 99.2 95.2 97.0 90.1 92.0 90.4 86.3
CLIP [28] 336 87.7 67.0 58.0 37.1 86.2 84.0 72.8 57.0 67.4 65.7 74.4 62.2
     +LLM2CLIP-15M 336 91.2 82.1 65.5 53.6 98.1 98.4 90.3 93.2 87.7 89.0 86.6 83.3
     +LLM2CLIP-60M 336 93.9 84.3 68.5 54.8 98.9 99.1 96.6 97.9 89.6 91.6 91.1 84.5
EVA02 [31] 336 89.6 78.0 64.2 47.9 91.5 89.4 76.6 70.0 74.7 76.4 79.3 72.3
     +LLM2CLIP-15M 336 94.7 85.1 69.7 58.5 98.8 99.2 93.5 96.2 89.2 91.3 88.0 84.8
     +LLM2CLIP-60M 336 95.9 85.4 72.4 58.8 98.8 99.2 95.5 97.3 90.7 92.6 91.0 86.8
SigLIP [38] 384 93.7 81.4 72.0 53.9 90.1 90.4 76.3 74.4 77.8 79.5 82.0 75.9
SO/14(428M)
SigLIP [38] 224 91.0 75.2 69.8 51.8 56.1 50.0 31.5 29.1 42.3 44.8 58.1 50.2
SigLIP2 [33] 224 93.9 82.9 72.0 55.5 90.2 87.2 75.7 74.5 77.0 78.9 81.8 75.8
     +LLM2CLIP-60M 224 95.2 84.6 73.7 57.6 99.4 99.4 96.2 97.3 91.5 92.9 91.2 86.4
SigLIP [38] 384 94.3 83.0 72.4 54.3 91.6 89.4 74.5 73.6 74.5 75.8 81.5 75.2
InternVL(6B)
InternVL [4] 224 94.2 81.6 69.4 53.2 92.4 90.5 79.6 77.7 79.2 81.6 83.0 76.9
Table 2: Systematic comparison of model performance on multiple datasets. The light blue row indicate LLM2CLIP post trained version.
Models Flickr-CN [16] COCO-CN [20] XM3600 [32]
I2T T2I I2T T2I I2T T2I
CN-CLIP 80.2 68.0 63.4 64.0
EVA-L-224 4.4 0.9 2.6 1.0 14.0 8.0
     +LLM2CLIP 90.6 75.6 72.0 70.1 68.3 56.0
Siglip2 79.2 56.9 55.3 51.7 59.7 48.2
     +LLM2CLIP 90.0 76.1 70.8 70.2 69.1 56.3
Table 3: Multi-lingua retrieval results on different datasets.
Methods Imagenet
0-shot Linear
CLIP L/14-336 76.6 84.8
+LLM2CLIP 74.1 85.2
CLIP L/14 75.5 83.9
+LLM2CLIP 71.8 84.4
CLIP B16 70.2 79.9
+LLM2CLIP 67.3 80.6
Table 4: Imagenet 0-shot classification and linear probe results.
MODEL VQA Datasets Pope Metrics MM Benchmarks Seed Benchmarks
VQAv2 GQA VizWiz SQA-IMG TextVQA Random Adv. Popular MME MMBench MMBench-CN LlavaBench MMVet All IMG Video
Llava (Paper) 78.5 62.0 50.0 66.8 58.2 87.3 86.1 84.2 1510.7 64.3 58.3 65.4 31.1 58.6 66.1 37.3
Llava (Rep.) 79.04 62.86 50.57 67.97 57.48 87.7 84.85 86.3 1476.69 66.66 60.39 58.0 34.3 59.86 66.95 39.71
     +LLM2CLIP 79.80 63.15 52.37 69.92 58.35 88.55 82.76 87.75 1505.82 68.29 60.40 62.7 34.8 60.96 68.80 38.96
Table 5: Performance of Llava 1.5 benchmarks. The best results are highlighted in bold. We explored whether LLM2CLIP could enhance complex image understanding tasks by modifying Llava’s visual encoder.
Methods Flickr COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
bge-en-icl 89.1 78.5 58.8 49.7 95.0 95.7 77.9 87.7 73.5 79.4 78.9 78.2
LLM2Vec-Llama-3-8B 91.1 81.4 61.5 51.9 94.5 96.4 82.1 88.6 77.7 82.6 81.4 80.2
NV-Embed-v2 90.4 80.0 60.5 51.6 94.5 95.7 83.3 90.0 78.3 82.1 81.4 79.9
Qwen2.5-0.5B-CC 86.2 74.9 56.0 45.1 92.6 93.2 73.7 79.0 69.5 73.0 75.6 73.0
Llama-3.2-1B-CC 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Llama-3-8B-CC 90.4 80.7 62.7 51.9 96.5 96.2 84.2 89.5 83.3 86.4 83.4 80.9
DeepSeek-R1-Distill-Llama-8B-CC 91.7 80.9 62.2 51.9 96.7 96.3 84.3 88.3 82.5 85.2 83.5 80.5
Llama-3.1-8B-CC 92.2 81.5 63.5 52.3 97.1 96.2 86.5 89.3 84.7 85.9 84.8 81.0
Llama3.1-8B 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
Table 6: Ablation for using different text encoder. BGE-EN-ICL, LLM2Vec-Llama-3-8B, and NV-Embed-v2 are community-released text encoders derived from LLMs. The suffix ”-CC” indicates encoders that have undergone our caption contrastive fine-tuning method based on the respective original LLMs.
Methods(MLLM’s Caption Ratio) Flickr COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
Directly Finetune(50%) 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
       +LLM2CLIP (100%) 89.2 77.2 55.6 47.6 97.6 98.0 88.5 92.5 82.6 86.9 82.7 80.4
       +LLM2CLIP (75%) 92.0 81.3 61.5 52.0 97.7 97.8 85.1 92.2 84.0 87.7 84.1 82.2
       +LLM2CLIP (50%) 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
       +LLM2CLIP (25%) 91.1 81.8 61.8 53.0 97.3 97.2 83.4 90.6 83.9 86.9 83.5 81.9
       +LLM2CLIP (0%) 91.1 81.5 60.9 53.0 96.0 94.4 72.0 85.6 82.6 85.6 80.5 80.0
Table 7: Ablation study on the proportion of dense captions from different MLLMs. Here, we explore the impact of both genuine image captions and fake dense captions generated by MLLMs on the performance of  LLM2CLIP.
Stage 1 Stage 2 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
89.2 77.7 59.0 47.3 94.3 94.3 75.0 80.5 74.0 77.5 78.3 75.5
Linear(×$\times 1$) 88.5 77.5 59.8 48.4 95.1 94.5 77.2 84.4 75.3 78.7 79.2 76.7
Linear(×$\times 2$) 89.6 78.1 59.6 48.5 95.5 94.7 79.7 84.0 76.1 79.0 80.1 76.8
Linear(×$\times 4$) 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Linear(×$\times 4$) Linear(×$\times 4$) 89.9 78.8 59.7 48.9 95.1 95.6 79.9 84.9 78.0 80.4 80.5 77.7
Transformer(×$\times 1$) 88.6 78.7 59.9 48.6 95.8 95.1 79.5 84.5 77.0 79.8 80.2 77.3
Transformer(×$\times 1$) Transformer(×$\times 1$) 89.8 78.6 58.1 49.0 96.1 95.4 81.9 85.6 76.5 78.1 80.5 77.3
[…]
3 Methods
3.1 Stage 1: LLM Caption Contrastive Fine-tuning
Traditionally, LLM output layers serve as classification heads producing text tokens. Prior works, such as LLM2Vec [2] and NV-EMBED-v2 [17], have explored converting LLMs into universal text encoders by applying additional designs and fine-tuning. Inspired by these approaches, we adapt similar concepts to the novel scenario of cross-modal pretraining. Our key design considerations encompass model architecture, training methods, and training data, summarized as follows:
Model Architecture:
  • Sentence Token Representation: We map the output layer features of the LLM to represent whole sentences. We considered two strategies: using the [EOS] token, which can attend to the entire sentence during pretraining, or employing average pooling across all output tokens. Empirically, average pooling performed better and is our default setting.
  • Bidirectional Attention: Given that the generative capability is unnecessary for our encoder-focused application, we remove the attention mask from the LLM, enabling bidirectional textual relationship modeling for enhanced comprehension.
  • Fine-tuning Parameters: To efficiently activate stronger textual comprehension capabilities within the output features of LLMs, we apply parameter-efficient fine-tuning using Low-Rank Adaptation (LoRA).
  • Adaptor Design: Adding an adaptor following the LLM can potentially enhance feature separability. We experimented with two adaptor variants: (1) a Transformer-based adaptor composed of several latent cross-attention transformer layers following NV-EMBED-v2, and (2) a Linear adaptor consisting of multiple linear layers directly attached after extracting sentence tokens from LLM outputs. However, neither approach significantly boosted CLIP performance during embedding fine-tuning. Consequently, we opted not to use adaptors during the LLM caption contrastive fine-tuning phase for simplicity.
Training Methods:
  • Masked Next Token Prediction (MNTP): Inspired by masked language modeling in BERT, MNTP improves feature quality by masking and predicting specific tokens. Unlike BERT, predictions are always made at the position preceding the masked token, aligning with LLM’s next-token prediction convention. While effective in pure text tasks (as shown by LLM2Vec), MNTP alone or combined with contrastive learning underperformed relative to direct contrastive fine-tuning in multimodal scenarios.
  • Textual Contrastive Learning: Following SimCSE, we employ textual contrastive learning to enhance feature separability, pulling positive samples closer and pushing negative samples apart in feature space. Two variants are possible: unsupervised (using dropout augmentation) and supervised (leveraging semantic pairs from annotated data). We default to supervised contrastive learning due to superior performance. Specifically, we generate positive pairs using two distinct captions from the same image, framed by a system prompt: ”Given a caption, retrieve a similar relevant caption.”
Training Data: We conducted experiments with Dreamlip [41]’s caption dataset, naturally containing multiple captions per image, making it ideal for supervised textual contrastive learning. Additionally, to ensure the generalizability of LLM textual differentiation capabilities, we incorporated a 1.5M paired pure-text dataset curated by Echo Embeddings [29] into our contrastive training process. 89213
3.2 Stage 2: LLM2CLIP Post Fine-tuning
We aim to perform post fine-tuning on pretrained CLIP since we believe the fundamental visual-language mapping has been well-established through CLIP’s extensive pretraining. By introducing an LLM, we expect to further enrich this representation space, potentially improving CLIP’s capabilities through a highly cost-effective post-training step. CLIP inherently consists of a Vision Encoder (typically a ViT) and a relatively small Text Encoder (roughly one-third the parameters of the Vision Encoder), structured as a compact autoregressive model. Integrating a caption-contrastive fine-tuned LLM as an additional text encoder and determining the optimal approach to enhance CLIP performance is critical. Here, we elaborate on model architecture, training methods, and training data.
Model Architecture:
  • Fine-tuning Parameters and Adaptor Design: Training CLIP typically requires large-scale datasets and significant batch sizes, resulting in considerable computational costs. Therefore, we aim to employ efficient Parameter-Efficient Fine-Tuning (PEFT) methods for this post-training stage. By default, we enable full gradients for the visual encoder to facilitate learning from LLM knowledge. Given the large size of the LLM, we compare two fine-tuning strategies: (1) directly fine-tuning the LLM using LoRA, and (2) freezing the LLM gradients and appending previously introduced Transformer or Linear adaptors after its output layer as learnable modules to obtain the sentence token features. Method (2) offers distinct advantages, including completely eliminating gradient updates for the LLM, thus significantly reducing GPU memory consumption. Moreover, since the adaptor is placed after the LLM, we can even precompute and store text features through offline inference, thereby reducing the inference overhead from multiple epochs to just a single pass. Consequently, this approach eliminates the need to load the LLM into GPU memory during training, greatly accelerating experiment iteration and training time, and drastically reducing GPU memory footprint. By default, we adopt method (2) with an adaptor consisting of four linear blocks. Detailed efficiency analysis will be presented in Section 4.4.
Training Methods: We tested several approaches:
  • (a) Removing the original CLIP Text Encoder and using LLM with CLIP Vision Encoder directly for cross-modal contrastive learning.
  • (b) Retaining both LLM and original Text Encoder, applying separate CLIP losses to both encoders.
  • (c) Adding an additional contrastive learning task between LLM and the original Text Encoder to method (b).
  • (d) Concatenating LLM and Text Encoder output features in method (b), then performing contrastive learning with CLIP Vision Encoder.
Extensive experiments (see Table 10) led us to adopt method (a) for its simplicity and effectiveness.
Training Data: With the integration of LLM, our method inherently excels at processing dense and detailed image captions. Datasets such as DreamLIP and Recaption, which leverage Share-Captioner [3], InstructBLIP [5] and LLaVA-1.5 [23] to generate multiple dense captions per image, can be effectively utilized. We explored multiple caption scenarios, employing multiple positive caption examples during contrastive training and blending real and MLLM-generated dense captions. Ultimately, we adopted a strategy blending real captions and dense MLLM captions at a 50% ratio.
Stage 1 Flickr COCO ShareGPT4V Urban-1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
Lora, AvgPool, Bidirectional, Supervise Simcse 88.9 78.8 59.8 49.1 96.3 95.6 80.1 85.1 77.0 80.8 80.4 77.9
Lora, AvgPool, Bidirectional, Un-supervise Simcse 77.2 66.0 48.8 38.2 80.2 86.2 47.7 53.5 41.9 44.7 59.2 57.7
Lora, AvgPool, Bidirectional, MNTP 82.0 71.7 51.2 43.1 92.0 91.9 67.0 70.3 58.4 58.2 70.1 67.0
Lora, AvgPool, Bidirectional, MNTP, Supervise Simcse 88.2 78.3 59.6 48.9 95.5 95.1 79.7 86.1 75.3 77.7 79.7 77.2
Lora, AvgPool, Casual, Supervise Simcse 89.4 78.7 59.9 49.0 95.4 95.7 79.9 84.9 75.3 79.1 80.0 77.5
Lora, EOS, Bidirectional, Supervise Simcse 89.6 78.6 59.4 48.5 95.3 95.4 79.8 83.8 76.0 80.0 80.0 77.3
Frozen, Linear Adaptor, AvgPool, Bidirectional, Supervise Simcse 85.5 75.5 54.1 45.3 92.0 93.5 72.3 75.5 66.4 66.9 74.1 71.3
Table 9: Ablation study on the training methods of LLM caption contrastive finetuning in Stage 1. Rows with a light blue background represent our default setting, and red text indicates content that differs from the default setting.
Methods Training Loss Testing Text Encoder Flickr COCO ShareGPT4V Urban1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP CL(CLIP-T, CLIP-V) CLIP-T 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
Directly Finetune CL(CLIP-T, CLIP-V) CLIP-T 89.3 77.8 59.3 48.5 88.3 88.2 68.5 76.0 67.2 71.2 74.5 72.3
   +LLM2CLIP a) CL(LLM, CLIP) LLM 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
b) CL(LLM, CLIP-V)+ CL(CLIP-T, CLIP-V) CLIP-T 89.6 77.8 59.4 48.6 88.1 87.7 68.0 74.8 67.0 71.3 74.4 72.0
LLM 90.1 81.3 61.3 52.4 97.1 97.4 85.6 91.1 83.8 86.9 83.6 81.8
c) CL(LLM, CLIP-V)+CL(CLIP-T, CLIP-V)+ CL(CLIP-T, LLM) CLIP-T 89.0 77.9 58.5 47.7 89.1 88.8 67.0 75.5 66.3 70.7 74.0 72.1
LLM 90.8 80.5 61.5 51.8 97.1 97.7 85.9 90.3 83.4 86.7 83.7 81.4
d) CL(LLM, CLIP-V)+CL(CLIP-T, CLIP-V)+ CL(Cat(CLIP-T, LLM), CLIP-V) CLIP-T 89.6 77.9 59.1 48.1 89.4 87.5 69.1 72.5 67.0 70.6 74.8 71.3
LLM 90.7 81.4 61.3 52.3 97.5 97.9 85.5 92.5 83.9 87.5 83.8 82.3
Cat(CLIP-T,LLM) 91.3 82.4 62.9 53.3 97.8 97.4 86.6 92.9 84.9 88.2 84.7 82.8
Table 10: LLM2CLIP Training Method Analysis. We experimented with various possibilities for post-training between the LLM and the pretrained CLIP model’s Vision Encoder (CLIP-V) and Text Encoder (CLIP-T). In the figure, ”CL” denotes the contrastive learning loss, and ”Cat” stands for concatenation. Items a–d correspond to the training methods section in Section 3.2. In this experiment, we performed zero-shot image-text retrieval tests on each text encoder, including the new feature projection space formed after concatenation.
Model Strategy Training Hour Batchsize Flickr COCO ShareGPT4V Urban1k DOCCI Average
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I i2t t2i
Llama-3.1-8B LLM LoRA 17 768 90.0 80.9 60.9 50.6 97.5 96.9 88.9 90.8 82.0 84.3 83.9 80.7
LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 85.2 74.6 52.3 40.8 87.7 89.6 57.5 59.2 49.9 48.3 66.5 62.5
Llama-3.1-8B-CC LLM LoRA 17 704 90.5 81.8 61.5 52.4 98.1 98.2 91.0 92.8 85.7 87.4 85.4 82.5
LLM Frozen + Linear Adaptor 5.5 4096 91.9 81.7 62.0 52.7 97.2 97.6 84.6 91.3 83.7 87.3 83.9 82.1
LLM Frozen + Linear Adaptor + Offline-loading 1.3 16384 93.2 82.2 64.2 54.0 98.2 98.3 87.8 93.1 86.2 88.8 85.9 83.3
Table 11: Efficiency study of LLM2CLIP. We use two nodes, each equipped with 8 A100 40GB GPUs, and conduct experiments by maximizing GPU memory utilization to achieve the largest possible batch size. We report training hours, batch size, and retrieval accuracy for each experiment. During offloading, we precompute text captions offline into embeddings using the LLM. This approach is a one-time preprocessing step that effectively reduces the experimental overhead.
4 Experiments
4.1 Default Experimental Settings
Default Settings for Caption Contrastive Fine-tuning We choose to use LLaMA 3.1 8B as the LLM integrated into CLIP pre-training, removing the causal attention mask to enable bidirectional attention. No additional adapters are used in caption contrastive finetune, and average pooling is employed to obtain the sentence token embedding. MNTP is not used by default.
We utilize 30M DreamLIP caption data to perform caption contrastive fine-tuning on the LLM. Randomly sampling any two captions of an image from DreamLIP as positive pairs, we conduct supervised SimCSE training.
Default Settings for LLM2CLIP Cross-modal Post-training After caption contrastive fine-tuning, we replace the original CLIP text encoder with the LLaMA 3.1 8B model to perform cross-modal contrastive learning post-training with the vision encoder. By default, we open the ViT gradients and freeze the LLM gradients, then add an adapter layer composed of four linear layers after the LLM. The design of the Adaptor is the same as FuseMix [35], using a simple inverted bottleneck MLP architecture. By default, we use 15M DreamLIP annotated subsets of CC 3M and CC 12M. For the 60M data setting, we use DreamLIP re-annotated CC 3M and CC 12M, YFCC 15M, and a 30M subset of the LAION dataset. For the 3M setting, we only use the CC 3M subset. We summerize our default experiments setting and implementation details in Appendix A.
4.2 System Comparison
In the system comparison, we used the default settings from  Section 4.1 throughout. However, we also conducted experiments on different pretrained SOTA CLIP models to verify that applying LLM2CLIP post training can significantly enhance their performance.
4.2.1 Zero-Shot English Text-Image Retrieval
Evaluation Dataset. For short-text retrieval, we used the MSCOCO [22] 5K test set and the Flickr [37] 1K test set. For long-text retrieval, we employed datasets from LongCLIP, including a 1K subset of ShareGPT4V [3], the Urban1K [39] dataset, and the DOCCI [26] dataset.
The ShareGPT4V-1M dataset consists of captions generated using GPT-4V and ShareCaptioner, covering images from LAION, CC, SBU [27], and MS COCO. Urban1K includes captions for 1,000 urban scene images, each richly annotated with detailed descriptions. DOCCI contains 1.5K high-resolution images with human-annotated captions and was used for retrieval evaluation. To facilitate observation, we additionally report the average Top-1 image-to-text (I2T) and text-to-image (T2I) accuracy across these five datasets.
LLM2CLIP Improves CLIP for Long- and Short-Text Retrieval. In Table Table 2, we compared OpenAI’s CLIP, EVA02, and Siglip2, and conducted experiments using three types of visual models: ViT-B/16, ViT-L/14, and SoVit/14. Regardless of the model size or resolution shown in the table, LLM2CLIP yields a significant performance boost for both CLIP and EVA02. For example, under EVA02 at 224 resolution, the LLM2CLIP-60M design achieves an average Top-1 retrieval accuracy improvement of +12 and +14.8. Even for SigLIP2—which was pretrained on 40B data—our method still brings an average performance improvement of +9.4 and +10.6; even on the short-caption datasets Flickr and COCO, where SigLIP2 is already very familiar, improvements of +1 and +1.9 are observed. Comparatively, the improvements for long captions are even larger, as the inherent large window of LLMs fully leverages their advantage in understanding long and complex texts.
Training Data Volume Analysis for LLM2CLIP. In Table Table 2, we also conducted experiments with 3M, 15M, and 60M training data volumes on CLIP and EVA02 using ViT-L/14. We found that, whether at 224 or 336 resolution, more data consistently leads to better retrieval performance, for both long and short text retrieval tasks. Furthermore, we observed that the 3M data experiment on EVA02 is already sufficient to yield a noticeable improvement in long text retrieval tasks; however, no clear gain was observed in short text retrieval. We speculate that with a small data volume, the model first needs to compensate for the loss incurred by disrupting the original cross-modal space—a fact evidenced by the significant difference between the 3M and 15M data experiments.
4.2.2 Cross-Lingual Retrieval
LLMs naturally enable strong multilingual capabilities. As in Table 3, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLip2.
4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
Evaluation Dataset. We conduct zero-shot retrieval testing on Imagenet, following the CLIP method. We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
Imagenet Performance Better In Linear Probe. Our method shows a slight drop in performance on Imagenet for zero-shot classification. We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions. However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement. This demonstrates that LLM2CLIP can indeed provide better visual features. One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
4.2.4 Multimodal Large Language Models Performance
Evaluation Dataset. Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
LLM2CLIP Improves VLLM Performance. As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
Experimental Settings. For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section. We use a subset of the CC 3M dataset for caption contrastive fine-tuning. Additionally, we explore augmenting the training data with the Wikitext-103 dataset [25] for MNTP training and the E5 dataset [29] as supplementary pure-text data for caption contrastive fine-tuning.
Structural Design Ablation. We examine various adaptor structures in Table 8. Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP post-training. Therefore, we default to using no adaptor structure in Stage 1.
Training Method Ablation. We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in Table Table 9, observing the following key insights: 1. Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential. 2. SimCSE is identified as the most critical loss function. MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance. 3. Supervised SimCSE substantially outperforms unsupervised SimCSE. 4. The performance difference between causal and bidirectional attention mechanisms is negligible.
Ablation for different LLMs. In Table 6, our method achieves significant improvements over these encoders. The result highlighted in red shows that using the original Llama3.1-8B without CC fine-tuning yields notably poor performance, negatively affecting the original CLIP.
4.4 Ablation Study: Stage-2 LLM2CLIP
Adaptor Design. As shown in Table 8, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training. Comparing Linear (×$\times 4$) and Transformer (×$\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure. Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance. Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
Training Method Ablation. We examine several methods of integrating LLMs into CLIP post-training in Table Table 10, with implementation details described in Appendix A. By comparing experiments (a)-(b) and (b)-(c), we find that neither jointly training two losses nor aligning text embeddings significantly improves performance. However, experiment (d), which applies contrastive learning on the concatenated encoder outputs, achieves notable improvements. For simplicity, we do not adopt this design by default but suggest it as a promising direction for future work.
Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP. As shown in  Table 11, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware. Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
Data Ratio Ablation. In Table Table 7, we experiment with various dataset compositions. The results show a clear trend: higher proportions of dense captions enhance performance in long-text image retrieval tasks, whereas lower dense caption ratios improve short-text retrieval performance. We select a 1:1 ratio by default, as it provides the most balanced improvement for both short and long texts.
Interestingly, even training exclusively on real short captions (100%), our model still significantly boosts performance for both short and long-text retrieval tasks. This improvement can be attributed to the incorporation of LLM, given that real captions share the same distribution as the CLIP pre-training data, yet LLM integration still yields substantial gains.
5 Conclusion
Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation, allowing language—the carrier of human knowledge—to play a pivotal role in disseminating the intelligence of LLMs across multiple modalities. This paper successfully proposes an efficient method to seamlessly integrate LLMs into CLIP post-training, demonstrating substantial performance improvements on state-of-the-art multimodal models, including CLIP, EVA02, and SigLip2.
Appendix A Implementation details
Details for Caption Contrastive Fine-tuning
We default to using LoRA for training the LLM, with a rank of 16 and $\alpha=32$. Training is conducted on four nodes, each with 8 NVIDIA A100 40GB GPUs, with an effective batch size of 2048 for one epoch, using a maximum sequence length of 512 tokens for 68 hours. The Adam optimizer is used with a learning rate of $2.5\times 10^{-4}$ and a linear learning rate warm-up for the first 300 steps. We employ bfloat16 quantization, gradient checkpointing, and FlashAttention 2 to optimize GPU memory consumption.
Details for LLM2CLIP Cross-modal Post-training
We employ the AdamW optimizer with a total batch size of 4096 and set the learning rate to $1\times 10^{-5}$ with a cosine decay rate of 0.05.
Details of multi encoder.
When using multiple encoders, we need to align CLIP-V, CLIP-T, and LLM embeddings. We design a one-layer linear projector for each encoder, maintaining the dimensionality at 1024. For the concatenation head scenario, we design a 2048-dimensional head placed after the concatenation of (LLM, CLIP-T) embeddings and another after CLIP-V.
Appendix B Additional results
Training Data Ablation.
In Table Table 12, we explore the impact of different data sources and volumes on Stage-1 LLM caption contrastive fine-tuning. Our findings include: 1. Comparing the combination of 3M Set1 and 1.5M E5 with 4.5M Set1 alone, adding supervised pure-text data from E5 does not provide practical benefits to CLIP’s cross-modal post-training performance and may even slightly degrade performance. 2. For datasets sampled differently from the same distribution (Set1 and Set2), the best results are achieved when Stage 1 and Stage 2 utilize distinct datasets. The overall retrieval performance may depend primarily on the total amount of data seen across both stages; for instance, the performance using Set2 in Stage 1 closely matches that of combining Set1 and Set2.
XM3600 Multilingual Results.
Figure 4 plots radar charts for zero‑shot T2I and I2T retrieval across 36 languages on the XM3600 dataset. Although Siglip2 is already pre‑trained on a large, multilingual corpus, augmenting it with LLM2CLIP—leveraging a large language model to enhance linguistic understanding—still yields a substantial boost in cross‑lingual retrieval performance, even when the fine‑tuning is performed exclusively on English data.